OpenStudy (anonymous):

The ratio of the present ages of two museums is 3 to 5. Twelve years ago, the sum of their ages was less than 56 years. Find the possibilities for the present age of the younger museum.

OpenStudy (anonymous):

Let the present ages of two museums be 3x and 5x Since, twelve years ago, the sum of their ages was less than 56 years !2 years ago age of first museum = 3x-12 and that of second museum = 5x-12 according to question, we have 3x-12 + 5x-12 =56 i.e. 8x-24 =56 i.e. 8x =56+24 i.e. 8x= 80 i.e. x= 10 Thus present age of first museum = 3x = 3*10= 30 years present age of second museum = 5x = 5*10= 50 years Hence present ages of both the museums is 30 years and 50 years.

OpenStudy (anonymous):

So the age of the younger museum is 30 years.

OpenStudy (anonymous):

so whats the answer???